OpenStudy (anonymous):

factor completely: 25s^2+4-20s

OpenStudy (mohsen):

(5x-2)(5x-2) x=2/5

OpenStudy (mohsen):

sorry,not x but s

OpenStudy (mathteacher1729):

As a general technique for solving these problems, we know we want our answer to look like ( stuff ) * (stuff) , to be more specific, we want it to look like ( number* s plus/minus number) * ( number*s plus/minus number). In this case you can see that the 25 multiplied by s^2 can factor only into 5*5 or 1*25. (that's where the 5x comes from). From there it's filling in the steps of FOIL.

OpenStudy (anonymous):

hold the phone. it says "factor" not solve for x. you do not have an equation here so do not write x = anything!

OpenStudy (mathteacher1729):

I would recommend checking out these videos: http://www.mathtv.com go to algebra > quadratic equation > factoring polynomials, and watch all the ones with the old guy. He is awesome. :D

OpenStudy (mohsen):

the value of s just an extra...

OpenStudy (anonymous):

it would help to write as \[25s^2-20s+4\] and recognize this as a "perfect square"

OpenStudy (anonymous):

\[25s^2\] is the square of \[5s\] 4 is the square of -2 and \[-20x =2\times -2\times 5x\]

OpenStudy (anonymous):

so right away you get \[(5x-2)^2\] because \[(a-b)^2=a^2-2ab+b^2\]

OpenStudy (mathteacher1729):

There is a nice illustration of that here: http://en.wikipedia.org/wiki/File:A_plus_b_au_carre.svg

OpenStudy (mathteacher1729):

Oh for goodness sake... http://upload.wikimedia.org/wikipedia/commons/c/c2/A_plus_b_au_carre.svg